OpenStudy (anonymous):

Two students are on a balcony a distance h above the street. One student throws a ball vertically downward at a speed vi; at the same time, the other student throws a ball vertically upward at the same speed. Answer the following symbolically in terms of vi, g, h, and t. (Take upward to be the positive direction.) how far apart are the balls at a time t after they are thrown and before they strike the ground?

OpenStudy (anonymous):

D = h+vi*t - ½g*t² - [h - vi*t - ½g*t²] = 2*vi*t

OpenStudy (anonymous):

@pully u could have drawn a diagram to make it better it would have been convincing...atleast for me